OpenStudy (anonymous):

sam bought brushes for $8,a palette for $5, and oil paints for $15. He paid $29.82 in all. What sales tax rate did sam pay?

OpenStudy (jdoe0001):

it says brushe(s), and paint(s), in plural, but looks he only got one maybe they had a get 2 for the price of 1 dunno :P

OpenStudy (jdoe0001):

$$ \text{n% of m is } \cfrac{n}{100}\times m\\ \text{so we know that}\\ \pmatrix{\cfrac{n}{100}\times m} + m = 29.82\\ \text{now we know m = (8+5+15) = 28, so}\\ \pmatrix{\cfrac{n}{100}\times 28} + 28 = 29.82\\ \text{now just solve for "n"} $$

OpenStudy (dan815):